Tags:
create new tag
view all tags

Shift Log 20111214

1935

SB/GW loading optics and tuning beam after FONT shift. Resetting damping ring orbit.

20:00

SB/GW Dispersion measurement & correction. Background around 3700 with 20dB of attenuation. Dispersion beforehand x:10mm and y:negligible.

Restored skew quads after FONT turned them off.

After correction with skew quads on, dispersion < 1mm in both x and y. mOTR emittance measurement 55pm in vertical. Using new correction method. Measurement afterwards 53pm.

SET11DEC11_0001.DAT optics loaded - BX 2.5 BY 1 optics by GW/MW.

14_20_46 set file. BY1 optics - background around 8000.

2100

Back to BX 10 BY 10 optics, background ~ 3200. 21_01 set file Access to replace FF lens, align post-IP beam dump, check OTR alignment and cover it over to reduce backgrounds. Laser referenced at between -35690 and -35600 on manipulator.

2250

Beam back on. Background 3700 counts, 9e9 charge at IP. OTR images captured - single beam at 200ms exposure file saved 10_59. Beam size estimate 5sigma from 420 - 513 px vertically. AA estimates this is (93/5)*((3*5.4)/7.88) = 38um - PSF, beam is smaller. Need correct polariser setting to analyse vertical size. Horizontally estimate 300px = 5sigma, gives sigmax = 123um. OTR position file 20111214_23_20_pos.dat Filter placed in OTR beam image 20111214_23_35_filteron. Using OTR beam image to search for notch in screen. Notch screen reference -36230 manipulator, CV 0, CH -150, wire scanner ~ 9000. Manipulator moved to -35645 (middle of laser range), CV moved to 585. Timing overlap.

Voltage phase shifter seems to wrap around 3.5V. Set to 0.7V. Had to turn OTR screen to 0.6 degrees to attenuate OTR. APD doesn't produce a voltage greater than 200mV and the OTR was saturating it. Remember different setup now without the recollimating LW lens. Added more attenuators to laser beam so that both the laser beam and the OTR were around 80mV. Able to double this at 0.7V on phase shifter.

CH: -150 CV: 550 - 650 10um steps - nothing
CH: -150 CV: 450 - 750 5 um steps - nothing
CH: -150 CV: 400 - 800 80 steps - nothing
CH: -350 CV: 400 - 800 80 steps - nothing
CH: +50 CV: 400 - 800 80 steps - nothing
CH: +250 CV: 400 - 800 80 steps - nothing
CH: -550 CV: 400 - 800 80 steps - nothing
CH: -150 CV: 300 - 900 150 steps - nothing
CH: -150 CV: 500 - 650 50 steps - nothing

ZH10X to -0.18A from -0.48A and ZH1FF -1.8A from -1.2A CH: -150 CV: 500 - 650 scan - nothing

Reset ZH10X to -0.48A and ZH1FF to -1.2A. Check timing and alignment.

CH: -250 CV: 400 - 800 80 steps - nothing

Phase scan 0 - 8V CH: -250 CV: 585 - nothing

Angle changes - ZH10X -0.68A from -0.48A and ZH1FF -0.507A from -1.2

CH: -250 CV: 400 - 880 120 steps - nothing

Angle change - ZH10X -0.28A and ZH1FF -1.71A

CH: -250 CV: 400 - 880 120 steps - nothing

ZH10X back to original setting -0.48A and ZH1FF to original setting -1.2A. Try scans further away horizontally.

CH: -1000 CV: 400 - 800 100 steps - nothing
CH: 600 CV: 400 - 800 100 steps - nothing
CH: 1200 CV: 750 - 800 25 steps - nothing
CH: 1200 CV: 400 - 800 80 steps - nothing

Access to check if any lens damage and to check laser energy transmission post IP.

Edit | Attach | Watch | Print version | History: r6 < r5 < r4 < r3 < r2 | Backlinks | Raw View | Raw edit | More topic actions

Physics WebpagesRHUL WebpagesCampus Connect • Royal Holloway, University of London, Egham, Surrey TW20 0EX; Tel/Fax +44 (0)1784 434455/437520

Topic revision: r6 - 11 Apr 2012 - LaurieNevay

 
This site is powered by the TWiki collaboration platform Powered by PerlCopyright © 2008-2023 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding RHUL Physics Department TWiki? Send feedback